# Unveiling galaxy chemical enrichment mechanisms out to cosmic dawn from direct determination of O & Ar abundances from JWST/NIRSPEC spectroscopy

### v1

@@ Appendix B Impact of line-blending on Ar abundances
Appendix C Impact of ionisation correction factors on Ar abundances
For Ar, the ionisation states of Ar2+, Ar3+, and in smaller amounts Ar+ and Ar4+ are possible for an SFG. Hence the ICF correction to the observed ionic abundances is relevant for Ar abundance determination. Arellano-Córdova et al. (2024) found that different ICF schemes, including Amayo et al. (2021) used here, have worked equally well for Ar abundance determination. They found that for z0.1 SFGs with 12 + log(O/H)8.2, ICF correction from observing only [ArIII] 7136 Å may underestimate the Ar abundance by up to dex in their sample, but no such effect is seen when 12 + log(O/H)8.2. This implies that the resulting log(O/Ar) values would be higher than their actual intrinsic values. The ICF correction based on the observation of both [ArIII] and [ArIV] lines is more accurate instead at all 12 + log(O/H) values (see their Figure 3). Only one of our SFGs (COSMOS 20062) has 12 + log(O/H)8.2 with its Ar abundance value determined from the observed Ar2+ lines only. Its Ar abundance value may indeed be underestimated but even considering the maximum offset, its log(O/Ar) abundance would still be consistent with MW-like chemical enrichment, as it is the galaxy with the highest log(O/Ar) in our sample (see Table 2). Other SFGs in our sample have 12 + log(O/H)8.2 and thus no ICF correction bias is expected.

### v2

@@ Appendix B Impact of line-blending on Argon abundances
Appendix C Impact of ionisation correction factors on Argon abundances
For Ar, the ionisation states of Ar2+, Ar3+, and in smaller amounts Ar+ and Ar4+ are possible for an SFG. Hence the ICF correction to the observed ionic abundances is relevant for Ar abundance determination. Arellano-Córdova et al. (2024) found that different ICF schemes, including Amayo et al. (2021) used here, have worked equally well for Ar abundance determination. They found that for z0.1 SFGs with 12 + log(O/H)8.2, ICF correction from observing only [ArIII] 7136 Å may underestimate the Ar abundance by up to dex in their sample, but no such effect is seen when 12 + log(O/H)8.2. This implies that the resulting log(O/Ar) values would be higher than their actual intrinsic values. The ICF correction based on the observation of both [ArIII] and [ArIV] lines is more accurate instead at all 12 + log(O/H) values (see their Figure 3). Only one of our SFGs (COSMOS 20062) has 12 + log(O/H)8.2 with its Ar abundance value determined from the observed Ar2+ lines only. Its Ar abundance value may indeed be underestimated but even considering the maximum offset, its log(O/Ar) abundance would still be consistent with MW-like chemical enrichment, as it is the galaxy with the highest log(O/Ar) in our sample (see Table 2). Other SFGs with [ArIII] 7136 Å detection in our sample have 12 + log(O/H)8.2 and thus no ICF correction bias is expected.
For ERO 10612, we have not utilised any ICF correction as no literature ICF scheme directly provides a recipe for Ar abundance determination when only Ar3+ ionic abundance is determined. Even if such a scheme were to come to pass, the Ar abundance of ERO 10612 would only increase by a given value, leading to a reduction in log(O/Ar) by the same value, thereby still keeping the same diagonal offset in Figure 4 from the MW-like chemical enrichment sequence in the log(O/Ar) vs 12 + log(Ar/H) plane.